OpenStudy (anonymous):

How to write (√2 +i√2)³ in trigonometric form?

OpenStudy (perl):

first change the complex number to trig form

OpenStudy (anonymous):

how do I do that

OpenStudy (perl):

it helps to graph the complex number on a modified x y coordinate plane. the y axis becomes the ' i ' axis

OpenStudy (perl):

the x axis stands for real number numbers

OpenStudy (anonymous):

it's the square root that's confusing me

OpenStudy (perl):

If we are given the complex number $$ \Large a + ib \\ \Large \text{this is equal to}\\ \Large r(cos(\theta) + i sin(\theta))\\ \Large \text {where}~~ r =\sqrt{a^2+b^2} ~,~ \theta =\arctan (b/a) $$

OpenStudy (anonymous):

i got r=2 and theta= 45 or pi/4 so 2(cos(pi/4)+isin(pi/4))?

OpenStudy (anonymous):

so does that mean it's 8(cos(3pi/4)+isin(3pi/4))?

OpenStudy (anonymous):

@perl

OpenStudy (perl):

$$ \large (\sqrt{2}+i\sqrt{2})^3 \\ \large =( 2 \cos (\pi/4) + i \sin(\pi/4))^3 \large= 2^3 (cos(\pi/4 * 3 ) +i \sin(\pi/4*3)) $$

OpenStudy (perl):

by demoivre's theorem

OpenStudy (anonymous):

oh, okay. Thank you!